OpenStudy (anonymous):

12-4/5(x+15)=4 (PLEASE SHOW WORK)

OpenStudy (anonymous):

is the '4/5' a fraction or division?

jhonyy9 (jhonyy9):

12-4/5(x+15)=4 40-4(x+15)=0 40=4(x+15) 40=4x+60 -20=4x x=-20/4 x=-5